A box of mass 80kg is placed on a rock surface if the frictional force between the box's and the surface is 200N calculate the coefficient friction


1
Expert's answer
2021-02-18T18:38:49-0500
"F_{fr}=\mu mg,""\mu=\dfrac{F_{fr}}{mg}=\dfrac{200\ N}{80\ kg\cdot9.8\ \dfrac{m}{s^2}}=0.25."